OpenStudy (anonymous):

According to the Americans with Disabilities Act, a ramp can rise no more than 1 ft for every 12 ft of horizontal distance. What is the maximum angle that a ramp can form with the ground?

OpenStudy (colouringme):

Are you familiar with sine cosine and tangent?

OpenStudy (anonymous):

yes but my teacher still says the answer i have is wrong i have 4.8 degrees as the answer is it wrong?

OpenStudy (colouringme):

wait

OpenStudy (colouringme):

yeah, can you write down how do you find the answer?

OpenStudy (anonymous):

tan(angle) = rise/run = 1/12 angle = atan(1/12) angle =~ 4.7636 degs --> 4.8 degs

OpenStudy (colouringme):

Are you sure of the value of atan? Make sure you've changed the mode on your calculator into degrees, because your steps are right, but the value of atan is wrong

OpenStudy (anonymous):

ohhh that makes sence

OpenStudy (colouringme):

Okay, keep me updated with your result! (:

OpenStudy (anonymous):

i got 3.7 this time and i switched the 2 numbers around

OpenStudy (colouringme):

Sorry, I think the first answer is correct, you have input the value of x and y correctly, really, really sorry ):

OpenStudy (colouringme):

maybe it should be 1.2 m, doesn't really make sense though